Find the generating function of recurrence relation an+1_an=3n ,n less than 0 where ao=1


1
Expert's answer
2022-02-01T17:26:20-0500

We have the recurrence relation "a_{n+1}-a_n=3n", for n less than 0, where "a_0=1".

Let's consider the sequence "b_m=a_{-m}", "m=-n", such that "m\in\mathbb{N}\cup\{0\}". Then

"a_{n+1}-a_n=a_{-m+1}-a_{-m}=b_{m-1}-b_m=3n=-3m", or "b_m-b_{m-1}=3m".

"b_m=(b_m-b_{m-1})+(b_{m-1}-b_{m-2})+\dots+(b_1-b_0)+b_0"

"=3m+3(m-1)+\dots+3+b_0=\frac{3}{2}m(m+1)+1"

The generating function of the sequence "a_n" is

"f(z)=\sum\limits_{n=-\infty}^0a_nz^n=\sum\limits_{m=0}^{+\infty}b_mz^{-m}="

"\sum\limits_{m=0}^{+\infty}\left(\frac{3}{2}m(m+1)+1\right)z^{-m}=\sum\limits_{m=0}^{+\infty}z^{-m}+\frac{3}{2}\sum\limits_{m=0}^{+\infty}m(m+1)z^{-m}"

"=\frac{1}{1-z^{-1}}+\frac{3}{2}z^2\frac{d^2}{dz^2}\sum\limits_{m=0}^{+\infty}z^{-m}=\frac{1}{1-z^{-1}}+\frac{3}{2}z^2\frac{d^2}{dz^2}\frac{1}{1-z^{-1}}"

"=\frac{1}{1-z^{-1}}+3\frac{z^2}{(z-1)^3}=\frac{1}{1-z^{-1}}+\frac{3z^{-1}}{(1-z^{-1})^3}"


Answer. The generating function of the sequence "a_n" is "f(z)=\frac{1}{1-z^{-1}}+\frac{3z^{-1}}{(1-z^{-1})^3}".